OpenStudy (anonymous):

what is the distance between (-4,-4) and (4,4)

OpenStudy (anonymous):

4sqrt(2)

OpenStudy (amistre64):

8 sqrt(2) seems more plausible

OpenStudy (amistre64):

i see an 8x8 right tri ...

OpenStudy (amistre64):

sqrt((4--4)^2 + (4--4)^2) sqrt(8^2 + 8^2) sqrt(64+64) = sqrt(2(64)) = sqrt(2).sqrt(64) = 8sqrt(2)